OpenStudy (anonymous):

r(t) | | | 1 | ________ | | | | | | |______|_______|_______________ 0 1 2 3 4 t I dont know how to read step functions.. How do I read this graph/find the equation of this graph????? PLEASE HELP

OpenStudy (anonymous):

Came out messed up. Point is, The switch starts at t=1 and ends at t=2.

OpenStudy (anonymous):

I want to know how to write out the equation for these graphs

OpenStudy (anonymous):

r(t)=u(t-1)-u(t-2)

OpenStudy (anonymous):

u is the unit step func.

OpenStudy (anonymous):

ooh i see i see.. and for example... let's say i wanted to graph 2t-2u(t-1).. would it become... (drawing it on paint)

OpenStudy (anonymous):

I think i've done the step function wrong... how would i do it?

OpenStudy (anonymous):

If I had a million guesses i'd never be able to do that... How come it goes up at a gradient like that?

OpenStudy (anonymous):

see u(t-1) starts at t=1 before t=1 your graph will be like y=2t...

OpenStudy (anonymous):

at t=1 r(t)=2*1-2=0 after t=1 ur u(t-1) has a const. value 1 so then ur equation r(t)=2(t)-2

OpenStudy (anonymous):

But why does the 2u(t-1) go up like that? I thought it was meant to be straight line :S Man that's confusing lol

OpenStudy (anonymous):

no...man ....u have a term 2t...that give u st. line...

OpenStudy (anonymous):

hello ggaborne...........

OpenStudy (anonymous):

Yep Thanks alot.. Helps!

OpenStudy (anonymous):

I have another question...

OpenStudy (anonymous):

I want to find the Inverse Laplace of \[(e ^{-t} - (e ^{-2t}) + (t^2))/(t)(t^2 -3t + 2)\] I was going to use partial fraction decomposition... but I can't get it working.

OpenStudy (anonymous):

now i'm really confused....but still it appears to me that partial fraction would work...

OpenStudy (anonymous):

Constantly getting the wrong answer... The question is... y'' - 3y' + 2y = r(t) y(0) = 1, y'(0) = 3 And r(t) is the graph i drew in the first post... The one where dipan said it is u(t-1) - u(t-2)

OpenStudy (anonymous):

how did u approached?

OpenStudy (anonymous):

I took the laplace of y'' - 3y' + 2y = u(t-1) - u(t-2) This gave me (t^2 - 3t + 2)Y(t) - t = (e^-t - e^-2t)/t Rearranging to get Y(t) on one side... I get the \[(e ^{-t} −(e ^{-2t})+(t^2))/(t)(t^2−3t+2)\] Then I get using partial.. \[[(e ^{-t} −(e ^{-2t})+(t^2))\] = A(t-1)(t-2) + B(t)(t-2) + C(t)(t-1) and i'm not sure how to find my unknowns A, B and C

OpenStudy (anonymous):

But even if i do find these constants.. A B and C.... it will turn out like the answer in the back of the book..

OpenStudy (anonymous):

hey...wait wait... u can't write [(e^−t−(e^−2t)+(t^2)) = A(t-1)(t-2) + B(t)(t-2) + C(t)(t-1) better to take this in the form of \[\frac{e^{-at}}{(t+b)^p}\]

OpenStudy (anonymous):

i'm confused...

OpenStudy (anonymous):

t(t^2-3t+2)=t(t-1)(t-2) \[\frac{e^{-t}-e^{-2t}+t^2}{t(t-1)(t-2)}=\frac{e^{-t}}{t(t-1)(t-2)}-\frac{e^{-2t}}{t(t-1)(t-2)}+\frac{t^2}{t(t-1)(t-2)}\]

OpenStudy (anonymous):

And then...? Find the partial decomposition of each one individually?

OpenStudy (anonymous):

yes but dont make any thing with the exponential terms...

OpenStudy (anonymous):

and \[\frac{e^{-at}}{(t+b)^p}\]gives u a direct form to solve it u can go to wiki for the details of that formula...

OpenStudy (anonymous):

Could you please give me the link from wiki? I don't know where I'd find it..